Use Euler’s method to solve the I.V.P.
y'=(x-y)/2, on [0,3], with y(0) = 1, and h = 0.5
1
Expert's answer
2020-04-21T16:51:40-0400

Given the initial value problem, "y'=\dfrac{x-y}{2}" with initial condition "y(0)=1".

The Euler's method states that,

"y_{n+1} = y_{n}+h\cdot f(x_{n},y_{n})", where "x_{n+1} = x_{n}+h".

Here, "h = 0.5, x_{0} = 0, y_{0} = 1, f(x,y) = y' = \dfrac{x - y}{2}" .

Step 1:

"x_{1} = x_{0} + h = 0.5\\
y_{1}=y(0.5)=y_{0}+h\cdot f(x_{0},y_{0})\\ = 1+ 0.5 \left(\dfrac{0-1}{2}\right) = 0.75"

Step 2:

"x_{2} = x_{1} + h = 1\\
y_{2}=y(1)=y_{1}+h\cdot f(x_{1},y_{1})\\ = 0.75+ 0.5 \left(\dfrac{0.5-0.75}{2}\right) = 0.6875"

Step 3:

"x_{3} = x_{2} + h = 1.5\\
y_{3}=y(1.5)=y_{2}+h\cdot f(x_{2},y_{2})\\ = 0.6875 + 0.5 \left(\dfrac{1-0.6875}{2}\right) = 0.765625"

Step 4:

"x_{4} = x_{3} + h = 2\\
y_{4}=y(2)=y_{3}+h\cdot f(x_{3},y_{3})\\ 
= 0.765625 + 0.5 \left(\dfrac{1.5-0.765625}{2}\right) = 0.94921875"

Step 5:

"x_{5} = x_{4} + h = 2.5\\
y_{5}=y(2.5)=y_{4}+h\cdot f(x_{4},y_{4})\\ 
= 0.94921875 + 0.5 \left(\dfrac{2-0.94921875}{2}\right)\\
= 1.2119140625"


Step 6:

"x_{6} = x_{5} + h = 3\\
y_{6}=y(3)=y_{5}+h\cdot f(x_{5},y_{5})\\ 
= 1.2119140625 + 0.5 \left(\dfrac{2.5-1.2119140625}{2}\right)\\
= 1.533935546875"